# Persistent and anti-Persistent Motion in Bounded and Unbounded Space: Resolution of the First-Passage Problem

### v1

@@ VII Iterative Procedure and Comparison Between Reflective Interactions
In the preceding sections we have exploited the analytical expressions of propagator generating functions. For the cases in which propagators are not known in closed form we present an iterative procedure to find the first-passage probability. While the procedure is valid in arbitrary dimensions, for notational ease we present it here explicitly for the one-dimensional case, and use it as a tool to make comparisons between the two forms of reflection mentioned in Sec. III.2. We differentiate between the two reflections based on whether an interaction with the boundary invokes a reversal in the direction of movement, which we call the ‘bouncing’ case, or if the movement direction persists, which we call the ‘bunching’ case (see panels (c) and (d) in Fig. 1 for pictorial comparison). For an in depth appreciation of the differences between the two boundary conditions we also compare here the occupation probability and the mean square displacement (MSD) in the two cases.
@@
where $\mathbb{D}$ represents the transition probabilities from state 1 to state 1, $\mathbb{E}$ from state 2 to state 1, $\mathbb{F}$ from state 1 to state 2 and $\mathbb{G}$ encoding the transitions from state 2 to state 2. Due to the finite space, each matrix is of size $N\times N$ and as each matrix only governs a subset of the movement possibilities, they are spars and have the following forms: $\mathbb{D}_{i,i}=c$, for $i\neq 1,N$, $\mathbb{D}_{i,i+1}=a$ and $\mathbb{D}_{1,1}=\delta+c$, $\mathbb{D}_{N,N}=\sigma+c$; $\mathbb{E}_{i,i+1}=b$ and $\mathbb{E}_{1,1}=\omega$; $\mathbb{F}_{i,i-1}=b$ and $\mathbb{F}_{N,N}=\omega$ and $\mathbb{G}_{i,i-1}=a$ and $\mathbb{G}_{1,1}=\sigma+c$, $\mathbb{G}_{N,N}=\delta+c$, where $\delta$, $\sigma$, $\omega$ are dependent on the chosen boundary conditions. With initial conditions $P_{n_{0}}(n,1,0)=\alpha_{1}$, $P_{n_{0}}(n,2,0)=\alpha_{2}$, that is $\bm{P}_{n_{0}}(1,0)=\alpha_{1}\bm{e}_{n_{0}}$ and $\bm{P}_{n_{0}}(2,0)=\alpha_{2}\bm{e}_{n_{0}}$, it is an iterative task to find the occupation probability at time $t$.
For the ‘bunched’ case, $\delta=b$, $\sigma=a$, $\omega=0$, while for the ‘bouncing’ case, $\delta=b$, $\omega=a$ and $\sigma=0$. To include absorbing traps one sets the outgoing probabilities of a trap state $(s,m_{s})$ as $(1-\rho_{s,m_{s}})A(s,m_{s},n,m)$ where $\rho_{s,m_{s}}$ is the partial absorption probability at $(s,m_{s})$ and with $A(s,m_{s},n,m)$ representing the elements of the transition matrix encoding movement out of $(s,m_{s})$. As an illustration, consider a trap at site $s=N$ and state $m_{s}=1$. In such a case we multiply the $N^{\text{th}}$ column of $\mathbb{D}$ and $\mathbb{F}$ by $(1-\rho_{N,1})$, and the first-passage is obtained via $F_{n_{0}}(n,t)=S_{n_{0}}(t-1)-S_{n_{0}}(t)$ with $F_{n_{0}}(n,0)=0$, where $S_{n_{0}}(t)=\sum_{n=1}^{N}P_{{n}_{0}}(n,t)$.
@@ VIII Conclusion
While there have been generalisations of renewal processes with continuous variables in the past, e.g., to non-Markovian dynamics [29] and ageing processes [82], they are limited by their analytic tractability, for example, obtaining only moments or asymptotic expansions. Our study instead generalises the renewal formulation in discrete space and time presented by Erdös, Feller and Pollard in 1949 and allows us to link analytically the occupation probability to the first-passage probability for persistent and anti-persistent walks with one or multiple targets. Our formalisms have highlighted the natural appearance of multiple modes in the first-passage dynamics when CRWs with strong persistence occur within finite periodic and reflecting domains. We have also quantified the directional first-passage statistics, that is the first-hitting dynamics based on the direction from which a walker reach a target.
@@ Appendix B The Structure Function
Presently, the matrix $\bm{\lambda}\left(\pi\mathcal{N}^{(\gamma)}_{\kappa_{1}},...,\pi\mathcal{N}^{(\gamma)}_{\kappa_{d}}\right)$ is of size $2d\times 2d$ and has a common structure for all the lattices we have considered here. Namely, the probability of remaining at the same state, i.e. persisting in the direction of the previous movement step or sojourning, is encoded in the diagonal elements, the probability of backtracking is given on the upper and lower diagonal in alternating rows, while all the remaining elements relate to the option of turning laterally. These features are evident in Eq. (4) for the $d=1$ case, and are seen explicitly also for the two-dimensional hypercubic lattice when we write out the matrix

### v2

@@ VII Iterative Procedure and Comparison Between Reflective Interactions
In the preceding sections we have exploited the analytical expressions of propagator generating functions. For the cases in which propagators are not known in closed form we present here an iterative procedure to find the first-passage probability. While the procedure is valid in arbitrary dimensions, for notational ease we present it here explicitly for the one-dimensional case, and use it as a tool to make comparisons between the two forms of reflection mentioned in Sec. III.2. We differentiate between the two reflections based on whether an interaction with the boundary invokes a reversal in the direction of movement, which we call the ‘bouncing’ case, or if the movement direction persists, which we call the ‘bunching’ case (see panels (c) and (d) in Fig. 2 for pictorial comparison). For an in depth appreciation of the differences between the two boundary conditions we also compare here the occupation probability and the mean square displacement (MSD) in the two cases.
@@
where $\mathbb{D}$ represents the transition probabilities from state 1 to state 1, $\mathbb{E}$ from state 2 to state 1, $\mathbb{F}$ from state 1 to state 2 and $\mathbb{G}$ encoding the transitions from state 2 to state 2. Due to the finite space, each matrix is of size $N\times N$ and as each matrix only governs a subset of the movement possibilities, they are sparse and have the following forms: $\mathbb{D}_{i,i}=c$, for $i\neq 1,N$, $\mathbb{D}_{i,i+1}=f$ and $\mathbb{D}_{1,1}=\delta+c$, $\mathbb{D}_{N,N}=\sigma+c$; $\mathbb{E}_{i,i+1}=b$ and $\mathbb{E}_{1,1}=\omega$; $\mathbb{F}_{i,i-1}=b$ and $\mathbb{F}_{N,N}=\omega$ and $\mathbb{G}_{i,i-1}=f$ and $\mathbb{G}_{1,1}=\sigma+c$, $\mathbb{G}_{N,N}=\delta+c$, where $\delta$, $\sigma$, $\omega$ are dependent on the chosen boundary conditions. With initial conditions $P_{n_{0}}(n,1,0)=\alpha_{1}$, $P_{n_{0}}(n,2,0)=\alpha_{2}$, that is $\bm{P}_{n_{0}}(1,0)=\alpha_{1}\bm{e}_{n_{0}}$ and $\bm{P}_{n_{0}}(2,0)=\alpha_{2}\bm{e}_{n_{0}}$, it is an iterative task to find the occupation probability at time $t$.
For the ‘bunched’ case, $\delta=b$, $\sigma=f$, $\omega=0$, while for the ‘bouncing’ case, $\delta=b$, $\omega=f$ and $\sigma=0$. To include absorbing traps one sets the outgoing probabilities of a trap state $(s,m_{s})$ as $(1-\rho_{s,m_{s}})A(s,m_{s},n,m)$ where $\rho_{s,m_{s}}$ is the partial absorption probability at $(s,m_{s})$ and with $A(s,m_{s},n,m)$ representing the elements of the transition matrix encoding movement out of $(s,m_{s})$. As an illustration, consider a trap at site $s=N$ and state $m_{s}=1$. In such a case we multiply the $N^{\text{th}}$ column of $\mathbb{D}$ and $\mathbb{F}$ by $(1-\rho_{N,1})$. The first-passage is obtained via $F_{n_{0}}(n,t)=S_{n_{0}}(t-1)-S_{n_{0}}(t)$ with $F_{n_{0}}(n,0)=0$, where $S_{n_{0}}(t)=\sum_{n=1}^{N}P_{{n}_{0}}(n,t)$.
@@ VIII Conclusion
While there have been generalisations of renewal processes with continuous variables in the past, e.g., to non-Markovian dynamics [29] and ageing processes [85], they are limited by their analytic tractability, for example, obtaining only moments or asymptotic expansions. Our study instead generalises the renewal formulation in discrete space and time presented by Erdös, Feller and Pollard in 1949 and allows us to link analytically the occupation probability to the first-passage probability for persistent and anti-persistent walks with one or multiple targets. Our formalisms have highlighted the natural appearance of multiple modes in the first-passage dynamics when CRWs with strong persistence occur within finite, periodic and reflecting domains. We have also quantified the directional first-passage statistics, that is the first-hitting dynamics based on the direction from which a walker reach a target.
@@ Acknowledgements
Data Availability
The data that support the findings of this study are available within the article and any supplementary files.
@@ Appendix A Reflective Boundary Conditions
Figure 11: (Colour Online). In the style of Fig. 1b in ref. [49], a schematic depicting the method to turn a periodic domain into a reflecting domain by squeezing the internal sites, while the boundary points (sites 1 and 4) remain unaffected by the transformation. For visual clarity we omit arrows depicting the state level dynamics, that is the directions from which a walker may enter a site, depicted in Fig. 2 of the main text. As shown in the right-hand panel, it is clear that after the transformation the walker may enter both internal states from either direction.
To employ the squeezing method, we transform a periodic domain into one with reflecting boundaries. The method, first introduced in ref. [49] for the standard LRW case can be applied also to the CRW and the essence of the method can be evinced from Fig. 11. As shown, if the walker occupies site 1 of the periodic lattice, its only options to move off the site are to travel to either site 2 or site 6. Thus, if we combine sites 2 and 6 together, that is we squeeze the lattice, to create a new site (labelled $2+6$ in the far right panel), we have precisely the reflecting boundary condition considered by Chandrasekhar [54].
@@ Appendix B The Structure Function
Presently, the matrix $\bm{\lambda}\left(\pi\mathcal{N}^{(\gamma)}_{\kappa_{1}},...,\pi\mathcal{N}^{(\gamma)}_{\kappa_{d}}\right)$ is of size $Z\times Z$ and has a common structure for all the lattices we have considered here. Namely, the probability of remaining at the same state, i.e. persisting in the direction of the previous movement step or sojourning, is encoded in the diagonal elements, the probability of backtracking is given on the upper and lower diagonal in alternating rows, while all the remaining elements relate to the option of turning laterally. These features are evident in Eq. (4) for the $d=1$ case, and are seen explicitly also for the two-dimensional hypercubic lattice when we write out the matrix